Tan square theta + Sec theta = 5 => Sec theta = 1) 32) 23) 14) -1

Grade:8

1 Answers

manna
7 years ago
 
Hi,
 
tan^2\theta +sec\theta =5
sec^2\theta -1+sec\theta -5=0
sec^2\theta +sec\theta -6=0
assume sec\theta =x
x^2+x-6=0
x^2+3x-2x-6=0
x(x+3) -2(x+3)=0
(x-2)(x+3)=0
x=2,x=-3 then x=sec thetha  put values in equation
sec thetha=2
hope i helped you...
 
and pls go through  trigonometry identities